The following sample observations were randomly selected


X 4 5 3 6 10 

Y 4 6 5 7 7 


a) Determine the correlation coefficient and interpret the relationship between X and Y. 

b) Determine the regression equation 

c) Determine the value of when X is 7. 

d) State two uses of goodness-of-fit test



1
Expert's answer
2021-08-03T11:17:42-0400

Solution:

Given:



(a):


"X\ Values:
\\∑ = 28
\\Mean = 5.6
\\∑(X - M_x)^2 = SS_x = 29.2

\\Y\ Values:
\\∑ = 29
\\Mean = 5.8
\\∑(Y - M_y)^2 = SS_y = 6.8

\\X\ and\ Y\ Combined:
\\N = 5
\\∑(X - M_x)(Y - M_y) = 10.6"


"r\ Calculation:
\\r = ∑((X - M_y)(Y - M_x)) / \sqrt{(SS_x)(SS_y)}

\\r = 10.6 / \sqrt{(29.2)(6.8)} = 0.7522"

So, the correlation coefficient=0.7522

Interpretation: This is a strong positive correlation, which means that high X variable scores go with high Y variable scores (and vice versa).

(b):

"\\Sum\ of\ products (SP) = 10.6

\\Regression\ Equation = ŷ = bX + a

\\b = SP/SS_X = 10.6/29.2 = 0.36301

\\a = M_Y - bM_X = 5.8 - (0.36\times5.6) = 3.76712

\\ŷ = 0.36301X + 3.76712"

(c):

Put x=7

"ŷ = 0.36301(7) + 3.76712
\\ŷ =6.30819"

(d):

Uses:

  1. The Chi-square goodness of fit test is a statistical hypothesis test used to determine whether a variable is likely to come from a specified distribution or not.
  2. It is often used to evaluate whether sample data is representative of the full population.